OpenStudy (anonymous):

it is given that 2 divided by 3 , 8 divided by d and n divided by 39 are equivalent fractions.find the value of d and n . :)

OpenStudy (anonymous):

2/3 = 8/d; 2d = 3*8; d = 24/2; d=12 2/3 = n/39; 3n = 2*39; n = 78/3; n = 26

OpenStudy (anonymous):

ok thnx